### v3

# Neural Algorithmic Reasoning with Multiple Correct Solutions

Abstract
Neural Algorithmic Reasoning (NAR) aims to optimize classical algorithms. However, canonical implementations of NAR train neural networks to return only a single solution, even when there are multiple correct solutions to a problem, such as single-source shortest paths. For some applications, it is desirable to recover more than one correct solution. To that end, we give the first method for NAR with multiple solutions. We demonstrate our method on two classical algorithms: Bellman-Ford (BF) and Depth-First Search (DFS), favouring deeper insight into two algorithms over a broader survey of algorithms. This method involves generating appropriate training data as well as sampling and validating solutions from model output. Each step of our method, which can serve as a framework for neural algorithmic reasoning beyond the tasks presented in this paper, might be of independent interest to the field and our results represent the first attempt at this task in the NAR literature.
1 Introduction
Classical algorithms such as Merge Sort guarantee correctness and generalize perfectly on abstract inputs [1]. Unfortunately, real problems are rarely abstract inputs. Consider pathfinding: to give traffic directions via a shortest paths algorithm, one must compress physical distance and multi-dimensional congestion data for each street into edge labels, thereby losing information. Bad edge labels mean bad directions, even despite a good algorithm. Neural Networks (NNs) avoid this problem by directly operating on high-dimensional data. Instead of compressing distance and congestion to a single path-length edge label, an NN can operate directly on distance and congestion data. The chief hurdle to NN performance is failure to generalize: common architectures are fragile to changes in input size and structure — exactly where algorithms succeed [1]. It is thus natural to wonder whether an NN might learn to generalize from an algorithm. To that end, Neural Algorithmic Reasoning trains NNs to mimic the running of classical algorithms [1]. A benefit of this approach is staying high-dimensional: no information is lost when projecting high-dimensional real data into a single abstract output for a classical algorithm [1]. Thus far, NNs have been trained to mimic algorithms which recover one solution from the space of possible correct solutions [2, 3, 4, 5, 6, 7, 8]. Ordinarily, an NN trained to perform DFS produces an array, which one can view as a distribution containing the probabilities of different solutions, from which the likeliest solution is returned as the model output. We adapt the model input to reflect multiple correct solutions and present methods of retrieving multiple solutions from the model output distribution. Using only one solution in NAR may be vulnerable to local minima in the model training. By diversifying the space of solutions returned by NAR, we may decrease the risk of solutions originating from a local minimum. For applications like cyberphysical systems, the flexibility of multiple solutions is essential. A single solution, for reasons of cost, safety, or compliance, may not be amenable to a final task.
1.1 Contributions
We give the first analysis of Neural Algorithmic Reasoners which return more than one solution. For training examples, we create distributions of solutions from multiple runs of classical algorithms with randomized tie-breaking. We train NNs to predict the generated distribution of solutions; and we leverage randomness in algorithm-specific ways to extract distinct solutions from the NN-predicted distributions. We also discuss methods of evaluating the solutions obtained through our methods. Thus we offer three contributions that might each be of independent interest: (1) a method for training a neural network on a distribution of correct solutions, (2) a method for extracting multiple solutions from the distribution the NN predicts, (3) an evaluation of the diversity and correctness of solutions our method yields. Together, these three contributions can serve as a framework for further research into Neural Algorithmic Reasoning with multiple solutions for each problem instance.
2 Methodology
To produce multiple solutions for a single graph, we train an NN to predict a distribution of solutions for each individual graph (solutions being individual DFS trees or BF paths, their distribution represented as distributions over child-parent pairs, as discussed in A.1.1 and A.2.1). Ordinarily, the CLRS benchmark [2] generates training data by running classical algorithms with deterministic tie-breaking. For example, the input graph in Figure 1 has two valid orders of depth-first traversal, but classical DFS will only return one based on which edge it explores first. To create our training examples, we run a classical algorithm (e.g. DFS) multiple times with randomized tiebreaking, and average the outputs (e.g. individual DFS trees) into a distribution of solutions. As we want the NN to predict a distribution, we train to minimize a distribution divergence (Kullback-Leibler: KL) [9] between the input distribution of solution and that given by the NN. Our final goal is to recover multiple solutions for an input graph; so, we develop stochastic, algorithm-specific sampling methods. Stochasticity of sampling methods means samples are likely to have multiple different solutions. For DFS, we sample the immediate parents of vertices until we reach the root vertex (which can be seen as executing DFS in reverse on a graph); our two DFS extractors, which we call Upwards and AltUpwards, differ in the manner in which they discard potential solutions (Appendix A.1.2). For BF, Beam samples a number of parents for non-source vertices according to parent-probabilities and chooses the lowest cost path arising (Appendix A.2.2). Also for BF, Greedy samples parents for each non-source vertex and chooses the parent with the lowest-cost outgoing edge for each non-source vertex (Appendix A.2.2).
Figure 1: Blueprint for NAR with multiple correct solutions. For Step 2 see Figure 3. For Step 4 see Figure 4 and Appendices A.1.2 A.2.2.
3 Experiments
Addressing the lack of evaluation methods for NAR with multiple solutions, we propose metrics to evaluate the validity and diversity of solutions. For a single graph, our NN predicts one solution distribution. From this distribution, we sample 5 solutions and count the number of distinct and valid solutions. For Bellman-Ford, a solution is valid if and only if it gives all shortest paths from the source vertex. For DFS, we say a solution is valid if and only if it does not contradict some necessary conditions A.2.3 (sufficient conditions not being known from tree structure alone). Figures 3 and 4 in Appendix A illustrate data generation and solution extraction on a graph. We investigate our coverage of the solution space in Appendix C. We run our experiments using the CLRS Benchmark [2], which handles data generation, model training, and model evaluation in a unified way. CLRS generates graphs of size 4, 7, 11, 13, and 16 to train our models on. In keeping with established methodology [2, 3], we test our models on graphs within and outside the range of our training sizes. For within-sample testing, we choose $n=|V(G)|=5$, as solutions on these graphs are easily manually verifiable and fast to compute at scale. For our out-of-distribution evaluation, we use the default CLRS-30 benchmark testing size $n=64$. The models are trained with default CLRS-30 benchmark model parameters (Appendix D) [2].
3.1 Results for Bellman-Ford
On small graphs, ($n=5$), sampling methods extract a single correct minimum-cost path with perfect accuracy from the NN-predicted distributions (Table 1, Table 2). To validate sampling methods, we also extract solutions from the empirical distributions. Similar results for empirical and predicted distributions suggest the NN predicts a good distribution for $n=5$ graphs. The lack of variety in solutions, with sampling methods always extracting the same solution, is expected for small graphs as there is likely only one minimum-cost path. Diversity increases in the case of $n=64$ graphs, where for each method, sampling a predecessor array five times gives a new solution in nearly all cases (Table 1). For larger graphs, Table 2 shows that sampling methods successfully extract correct minimum-cost paths from the empirical distributions (denoted as $\mathcal{P}$). However, extraction is often inaccurate from the distributions of solutions predicted by our models (denoted $\tilde{\mathcal{P}}$). These results permit us to conclude that our models perform well on small graphs, but have difficulty scaling outside of their training distribution. While comparing our approach with the deterministic algorithmic frameworks of Veličković et al. [2], Ibarz et al. [3] is disadvantageous to our method due to its broader scope, it is worth noting that our BF results are markedly worse than the 99% micro-F1 score in [2, 3]. However, our results are encouraging for future work, as we have demonstrated the feasibility of training neural networks to emulate a version of Bellman-Ford that is more powerful than its deterministic variant.
Table 1: Mean proportions of unique and valid solutions when sampling 5 solutions from empirical ($\mathcal{P}$) and model output ($\tilde{\mathcal{P}}$) distributions for each graph in 5 training runs. Rounded figures.
Bellman-Ford
Uniques ($\mathcal{P}$) Valids ($\mathcal{P}$) U ($\tilde{\mathcal{P}}$) V ($\tilde{\mathcal{P}}$)
Greedy ($n=5$) $0.20\pm 0.00$ $1.00\pm 0.00$ $0.20\pm 0.00$ $1.00\pm 0.00$
Beam ($n=5$) $0.20\pm 0.00$ $1.00\pm 0.00$ $0.20\pm 0.00$ $1.00\pm 0.00$
Greedy ($n=64$) $1.00\pm 0.00$ $1.00\pm 0.00$ $1.00\pm 0.01$ $0.40\pm 0.19$
Beam ($n=64$) $1.00\pm 0.00$ $1.00\pm 0.00$ $1.00\pm 0.00$ $0.54\pm 0.22$
Depth-First Search
Uniques ($\mathcal{P}$) Valids ($\mathcal{P}$) U ($\tilde{\mathcal{P}}$) V ($\tilde{\mathcal{P}}$)
Upwards ($n=5$) $0.74\pm 0.04$ $0.32\pm 0.06$ $0.52\pm 0.05$ $0.16\pm 0.07$
AltUpwards ($n=5$) $0.51\pm 0.03$ $0.86\pm 0.02$ $0.73\pm 0.05$ $0.56\pm 0.14$
Upwards ($n=64$) $1.00\pm 0.00$ $0.01\pm 0.02$ $1.00\pm 0.00$ $0.00\pm 0.01$
AltUpwards ($n=64$) $1.00\pm 0.00$ $0.05\pm 0.03$ $1.00\pm 0.00$ $0.01\pm 0.02$
3.2 Results for Depth-First Search
In most cases, our sampling methods perform worse than simply using the most likely predecessor for all $v\in V(G)$, but have the advantage of recovering multiple solutions (Table 1). On smaller graphs, AltUpwards performs well on the model distributions, returning a valid DFS-tree around half of the time. We only consider necessary conditions, causing up to 40% of trees identified as "correct" to be incorrect for $n=5$ (evaluating this metric for $n=64$ is computationally prohibitive in absence of sufficient correctness conditions). DFS is known to be difficult, with NARs giving correct (single) solutions only around a quarter of the time without pointer reversal [2, 3, 4]. This contextualises our results and demonstrates the promise of our approach compared to both classical algorithms and previous work in neural algorithmic reasoning. In terms of the ability of recovering multiple solutions, Table 1 shows that for both considered graph sizes, many unique solutions are obtained, with better accuracy scores for $\mathcal{P}$ than for $\tilde{\mathcal{P}}$, particularly for large graphs, where accuracy is low in general. In fact, for the out-of-distribution testing, our NAR method for DFS performs significantly worse than the current state of the art [3]. Because our method is non-deterministic, this should not be taken as a conclusively negative result: we would expect our method for returning multiple solutions to perform worse than the results for models giving single solutions in [3], as predicting multiple solutions is a more difficult task with more noise. Furthermore, AltUpwards is generally able to recover more unique and more valid solutions than Upwards, but at a high computational cost. Table 1 shows that for large graphs, both sampling methods yield many unique solutions, but these solutions are hardly ever valid for large graphs.
Table 2: Mean accuracies (0 to 1) for all sampling methods, with standard deviations over 5 runs.
Bellman-Ford
Argmax Beam Greedy Random
True (n = 5) $1.00\pm 0.00$ $1.00\pm 0.00$ $1.00\pm 0.00$ $00.00\pm 0$ 0.00
Model (n = 5) $1.00\pm 0.00$ $1.00\pm 0.00$ $1.00\pm 0.00$ $00.00\pm 0$ 0.00
True (n = 64) $1.00\pm 0.00$ $1.00\pm 0.00$ $1.00\pm 0.00$ $00.00\pm 0$ 0.00
Model (n = 64) $0.50\pm 0.28$ $0.56\pm 0.18$ $0.38\pm 0.19$ $00.00\pm 0$ 0.00
Depth-First Search
Argmax AltUpwards Upwards Random
True (n = 5) $0.80\pm 0.10$ $0.90\pm 0.07$ $0.34\pm 0.18$ $00.00\pm 0$ 0.00
Model (n = 5) $0.72\pm 0.28$ $0.54\pm 0.11$ $0.14\pm 0.15$ $00.00\pm 0$ 0.00
True (n = 64) $0.10\pm 0.07$ $00.00\pm 0$ 0.00 $0.04\pm 0.05$ $00.00\pm 0$ 0.00
Model (n = 64) $0.14\pm 0.21$ $0.04\pm 0.05$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
Conclusion
We outline a two-part approach to NAR with multiple solutions: we train models to predict distributions of solutions and stochastically extract solutions from those distributions. We demonstrate feasibility, given a good distribution, but also that NNs do not easily learn to predict good distributions. We attribute this difficulty to a number of factors. First, NAR is difficult: previous work has shown that NNs only achieve up to 40% micro-F1 on deterministic DFS without hint pointer reversal, which improves it to nearly 100% [4]. Second, training a model to fit a distribution of solutions does not directly optimize for correct solutions. Solutions must still be extracted from the distribution via stochastic methods, to which the loss metric does not extend (Figure 7). Improving methods for extracting solutions remains an important algorithmic task. Still, with the promise of sampling multiple solutions from a distribution demonstrated, we believe our modifications of the CLRS benchmark act as proof-of-concept for further work. Our method is applicable to other CLRS-30 algorithms such as Minimum Spanning Tree, Bipartite Matching, and Task Scheduling. Having multiple solutions adds essential flexibility to applications of NAR such as cyberphysical systems and might aid in explaining NAR models.
Appendix A Methodology
We discuss our methods for randomising the deterministic (canonical) versions of Depth-First Search and the Bellman-Ford algorithm in order to use the resulting distributions over solutions as inputs to the CLRS-30 baseline. We also present our methods of extracting candidate solutions from model outputs and give approaches towards verifying them as correct solutions of the respective deterministic algorithms our sampling methods are aiming to emulate.
Figure 2: Parent Tree encoding of a single DFS solution. The start node, 0, is its own parent, represented by the value 0 at index 0. The value 2 at index 1 indicates that vertex 2 is the parent of vertex 1. Similarly, the value 0 at index 2 indicates that vertex 0 is the parent of vertex 2.
Figure 3: Parent Probability Distribution encoding of multiple DFS solutions. The value 0.5 in row $1$ column $2$ indicates that vertex 2 is a parent of vertex 1 in 50% of correct parent trees. We call the parent probabilities obtained by repeatedly running an algorithm the true parent distribution $\mathcal{P}$. The model outputs, which we aim to be as close as possible to $\mathcal{P}$, are denoted as $\tilde{\mathcal{P}}$.
Figure 4: DFS tree extraction from a parent probability matrix: $\tilde{\mathcal{P}}$ or $\mathcal{P}$. The parent tree [0,2,2] is incorrect because there is no solution in which 2 is its own parent. [0,0,0] is incorrect, despite there being solutions in which 0 is the parent of each vertex, because there is no solution in which 0 is the parent of every vertex: if 0 is the parent of 1, 1 must be the parent of 2.
A.1 DFS
We consider a generalization of the deterministic DFS setting implemented in the CLRS benchmark [2]. In our setting, edge-exploration is random: from a given vertex, any of the outgoing edges may be explored first. Still, we require that once DFS backtracks at the root of a connected component, it restarts its search at the next vertex in a linear ordering of the vertices. We think the generalization is natural, modelling a setting where one conducts DFS according to a chosen sequence of starting points, without assumptions about which edges to pursue first.
Figure 5: A graph with multiple correct DFS solutions. Graph edges are drawn in black. The red tree would be the result if vertices were explored in alphabetical order, while the blue tree is one of multiple other posible valid DFS trees. Note also how the DFS-tree is encoded as a predecessor array (with the entry at position $i$ is the predecessor of the $i^{th}$ vertex in alphabetical order, which we henceforth call $\pi$.
A.1.1 Randomness in training
We generate a ground-truth parent distribution $\mathcal{P}$ for each graph in the test set, and train the model to minimize the KL-divergence between its predicted distribution $\tilde{\mathcal{P}}$ and $\mathcal{P}$. We obtain $\mathcal{P}$ by combining 20 solutions generated by a DFS algorithm that randomly tiebreaks between outgoing edges. Refer to Appendix B for a justification of using 20 solutions. We compute the fraction of solutions in which node $s$ is the parent of node $t$, which we say is the parent probability, as shown in Figure 3. Algorithm 1 gives detailed description of our particular DFS algorithm, with the line that randomizes edge-exploration highlighted in red.
Algorithm 1 Randomised Depth-First Search
1:Input: Adjacency Matrix $A$, vertices $V$ in linear order
2:Output: Parent Tree $\pi$, encoded as list: $\pi[i]:=$ parent of $i$
3:Initialize color $\leftarrow[0]*|V|$triangleright list of |V|-many 0s
4:Initialize startNode $\leftarrow V[0]$
5:rootOrdering $\leftarrow$ $V$triangleright ordered restarts
6:tiebreakOrder $\leftarrow$ shuffle$(V\setminus V[0])$triangleright random edge-exploration
7:for root in rootOrder do
8:     if color[root] == 0 then $\triangleright$ vertex not yet explored
9:         color[root] $\leftarrow 1$
10:         $\pi$[root] $\leftarrow$ root
11:         current $\leftarrow$ root $\triangleright$ pointer to node being explored
12:         $s_{last}\leftarrow$ root $\triangleright$ pointer to node most-recently found
13:         $s_{prev}[s_{last}]\leftarrow$ root $\triangleright$ pointer to node before the node most-recently found
14:         while True do
15:              for potentialChild in tiebreakOrder do
16:                  if A[root][potentialChild] != 0 then
17:                       if color[potentialChild == 0] then
18:                           color[potentialChild] $\leftarrow$ 1
19:                           $\pi$[potentialChild] $\leftarrow$ current
20:                           $s_{last}\leftarrow potentialChild$
21:                           break  $\triangleright$ child found, exit for-loop, continue while
22:                       end if
23:                  end if
24:              end for
25:              if current == $s_{last}$ then $\triangleright$ No outgoing edge found
26:                  if $s_{prev}[s_{last}]==s_{last}$ then $\triangleright$ Done: backtracked source and no new edges
27:                       break $\triangleright$ Exit while loop, try further with next root in rootOrder
28:                  end if
29:                  $s_{last}\leftarrow s_{prev}[current]$triangleright Prepare to Backtrack
30:              end if
31:              current $\leftarrow s_{last}$triangleright go-deeper if new node found, else backtrack
32:         end while
33:     end if
34:end for
A.1.2 Randomness in sampling
We compare 4 methods for extracting valid DFS solutions from a parent distribution, $\tilde{\mathcal{P}}$. First, argmax sampling assumes the parent of each node $x$ to be the node $u$ such that
$u=\underset{v\in V(G)}{\text{max}}\tilde{\mathcal{P}}_{v,u},$ (1)
where $\tilde{\mathcal{P}}_{v,u}$ is the probability of an edge $(v,u)$ according to $\tilde{\mathcal{P}}$. Argmax sampling is deterministic and can return only one solution. We include it in our discussion as a good baseline which corresponds to the internal mechanism of the CLRS-30 benchmark [2]. To extract multiple solutions for DFS, we introduce stochasticity according to Upwards Sampling, which is described in Algorithm 2.
Algorithm 2 Upwards Sampling
1:Input: $\tilde{\mathcal{P}}$
2:Output: A sampled predecessor array $\pi$
3:Initialize $\pi$
4:leaves $\leftarrow sort(\sum_{v\in V}\tilde{\mathcal{P}}_{u,v}\text{ for all }u\in V)$triangleright Sort vertices by their probability of being leafs
5:while $leaves\neq\emptyset$ do
6:     leaf $\leftarrow leaves[0]$triangleright Take the "leafiest" vertex
7:     $\pi[leaf]\leftarrow SamplePredecessor(\tilde{\mathcal{P}},leaf)$triangleright Sample $\pi[leaf]$ according to $\tilde{\mathcal{P}}_{leaf}$
8:     leaves $\leftarrow leaves.remove(leaf)$triangleright Prevent leaf from being considered as a child again
9:     for $v\in V$ do
10:         $\tilde{\mathcal{P}}_{leaf,v}\leftarrow 0$triangleright Prevent leaf from being sampled as a parent
11:     end for
12:     leaf $\leftarrow\pi[leaf]$triangleright Make predecessor the next leaf
13:     while $\pi[leaf]$ is undefined do $\triangleright$ Sample the predecessor’s predecessor
14:         $\pi[leaf]\leftarrow SamplePredecessor(\tilde{\mathcal{P}},leaf)$
15:         leaves $\leftarrow leaves.remove(leaf)$
16:         for $v\in V$ do
17:              $\tilde{\mathcal{P}}_{leaf,v}\leftarrow 0$
18:         end for
19:         leaf $\leftarrow\pi[leaf]$
20:     end while
21:end while
In Upwards Sampling (Algorithm 2), we first determine how likely each node is to be a parent node of any other node. We do so by summing the probabilities of each vertex $u\in V$ to be the parent of each vertex $v\in V$ and then sorting this list in ascending order, as done in Line 4. In each iteration of the outer loop (Line 5), we begin by taking the "leafiest" vertex (the vertex with the lowest probability of being a parent) and sampling its parent according to $\tilde{\mathcal{P}}_{leaf}$ (Line 7). After having done so, we remove the leaf from future consideration as a leaf and as a parent (as a leaf by definition will not be the parent of another node). Next, we sample the parent of $\pi[leaf]$ until we encounter a node that already has a parent (Line 5). This ensures we do not overwrite previously sampled parents. We have chosen this sampling method as it works similarly to DFS while respecting the inherent stochasticity resulting from returning $\tilde{\mathcal{P}}$ rather than one unambigous solution, as has been done in [2] and [3]. During testing, we encountered the situation that the removal of leaves from consideration as parents leads to no potential parents being available for a node due to having been removed. To address this problem, we give a node a random parent in this case. We further implement a method without this removal, which we call alt Upwards Sampling. While less faithful to graph properties, it addresses the limitations caused by the difficulty of computing the "leafiness" of each node while operating closer to DFS. It can be seen in Section 3.2 that this method outperforms Algorithm 2.
A.1.3 Checking Validity
Reconstructing multiple solutions from a matrix of probabilities introduces the possibility of error. We test the solutions we extract from our models against the following necessary conditions for correctness.
We say purported parent tree $\pi:=(v_{0},v_{1},...,v_{|V|-1})$ is DFS-valid for a graph $G=(V,E)$ with linearly ordered vertices $V$ only if the following 5 conditions hold:
  • The start node is correct:
    The least element of the linear ordering on Vertices is its own parent: $\pi[0]:=v_{0}=0=V[0]$.
  • The edges are correct:
    All edges suggested by the parent tree are in fact edges in the graph:
    $\forall t\in V,t==\pi[t]\lor E(\pi[t],t)$.
  • There is no cycle bigger than a self-loop in the parent tree:
    $\neg\exists C\subseteq V:\forall v\in C,\exists s,t\in C,s\neq v,t\neq v:E(s,v)\land E(v,t).$
  • Each vertex which is its own parent is not reachable by a lower-index node:
    $\forall t\in V,\pi[t]==t\implies\forall s\in V,s<t,\neg path(s,t)$.
  • Each parent whose child is not itself is reachable from the lowest index node for which its child is reachable: $\forall t\in V,\pi[t]\neq t\implies\exists l\in V:path(l,\pi[t])\land\forall s\in V,path(s,t)\implies l\leq s$111path(x,x) is true
No sufficient conditions to determine whether, given a tree $T$ and a graph $G$, $T$ is a valid DFS-tree of $G$ (for some tie-breaking rule) are known to the authors in the graph theory literature.
A.2 Bellman-Ford
The Bellman-Ford algorithm computes the minimum-cost paths from a single source vertex $s$ to each other vertex $v$ in the graph. As is the case for DFS, the solutions are represented as a predecessor array $\pi$.
A.2.1 Randomness in training
Similarly to randomising DFS, our implementation of a randomised Bellman-Ford algorithm relies on randomising the order in which paths are considered by shuffling the lists of nodes. Figure 6 illustrates a situation in which multiple valid Bellman-Ford predecessor arrays are possible. However, the final output of the canonical Bellman-Ford algorithm will only be influenced by such a randomisation procedure if there are multiple minimal-weight paths from the source node $s$ to at least one $v\in V\backslash\{s\}$. As the CLRS-30 benchmark samples Erdős-Rényi graphs with real-valued edge weights, the probability of two paths $(s,v)$ for any $v\in V\backslash\{s\}$ having the same cost is vanishingly low. To this end, we use a smaller set of possible weights such that for each weight $w_{ij}\in\{1,2,3\}$, which we normalise for training purposes. As for DFS, we obtain a distribution of parent nodes by re-running the algorithm 20 times and computing the frequency of each parent for each node.
Figure 6: A graph with multiple possible minimum-cost $(source,v)$-paths. As previously, the black arrows represent the edges, with the red and blue edges showing different minimum-cost paths corresponding to their respective predecessor arrays.
A.2.2 Randomness in sampling
We use an algorithm we call BF Beamsearch to sample predecessor arrays $\pi$ representing the minimum-cost $(s,v)$-paths for each $v\in V(G)$. This algorithm closely resembles standard Beamsearch, but it differs as it samples according to the distribution given by the model. For each $v\in V(G)$, BF Beamsearch considers paths from the source to $v$ of length at most $n$ by first sampling candidate predecessors of $v$ and then recursively sampling predecessors of predecessors until the path has length $n$. At each stage, only the 3 lowest-cost paths are kept and used for the next sampling. As soon as the maximum recursion depth is reached, the predecessor that is at the end of the lowest-cost path from the source is chosen as $\pi[v]$. This mimics the operation of the deterministic Bellman-Ford algorithm while respecting randomness and is thus a suitable, albeit involved, way of sampling $\pi$. To sample predecessor arrays more easily from $\mathcal{P}$ and $\tilde{\mathcal{P}}$, we use a greedy randomised algorithm. This algorithm samples a set of parents for each vertex $v$ except the source node (which we assume to be its own parent) according to $\tilde{\mathcal{P}}_{v}$. In the next step, it checks whether any of the sampled parents are plausible (i.e., if there exists an edge $(p,v)\in E(G)$ for the parent $p$). It resamples if there are no plausible parents. If any of the sampled parents are plausible, or if the maximum number of resamplings is reached, it chooses the plausible parent $p$ with the lowest weight $w_{p,v}$ as the predecessor $\pi[v]$. Furthermore, we also use the methods of deterministically choosing the likeliest parent of $v$ according to $\tilde{\mathcal{P}}_{v}$ and randomly uniformly choosing a parent as was done with DFS in order to obtain some baseline results.
A.2.3 Checking Validity
In contrast to DFS, the task of deciding whether a predecessor array sampled from a model output distribution $\tilde{\mathcal{P}}$ is a valid output of running the Bellman-Ford algorithm is straightforward. To this end, we implement a simple algorithm called check valid BF path, which is given in Algorithm 3.
Algorithm 3 check valid BF path
1:Input: Adjacency Matrix $A$, source vertex $s$, sampled predecessor array $\pi$
2:Output: Decide whether $\pi$ is a correct Bellman-Ford solution
3:$G^{\prime}\leftarrow buildGraph(\pi)$triangleright Construct a graph given by the paths in $\pi$
4:if $E(G^{\prime})\not\subseteq E(G)$ then $\triangleright$ Check whether $G^{\prime}$ contains any nonexistent edges
5:     Return False
6:else
7:     $true\_costs\leftarrow BellmanFord(A,s)$triangleright Compute the costs using the Bellman-Ford algorithm
8:     $model\_costs=PathCosts(\pi)$$\triangleright$ Compute costs of paths in $\pi$
9:     Return $model\_costs==true\_costs$
10:end if
Algorithm 3 simply constructs a graph from the vertices and edges that are implicitly given in $\pi$ and checks whether all edges correspond to edges in the graph $G$ and whether the costs of the paths in $\pi$ are the same of the costs when using the deterministic Bellman-Ford algorithm on $(G,s)$.
Appendix B Using different numbers of algorithm re-runs
When creating our training data by generating distributions over parents for DFS and Bellman-Ford, we need to consider how often we run the randomised versions of the deterministic algorithms on a graph in order to obtain a good distribution over solutions. To this end, we consider the graph sizes $n\in\{5,\dots,64\}$ and generate 100 graphs for each size and run an algorithm (DFS or BF) 20, 50, and 100 times for each graph. For each graph, we then consider the KL-Divergence between each pair of distributions generated by a different number of algorithm re-runs (we compare the distribution generated from 20 runs to the distribution generated from 50 runs and to the one generated from 100 runs, et cetera). We plot the mean KL-Divergence along with its standard deviation for each pairwise comparison between re-run numbers for all considered sizes. In Figure 7, we see that while there are differences between the distributions generated from a varying number of algorithm re-runs, the distribution differences do not change markedly when going from 20 to 100 versus from 50 to 100 re-runs. Therefore, by increasing the number of re-runs, it appears that we do not gain enough information to justify the added computational cost of an increased re-run number. Therefore, we settle on 20 as a reasonable value for the number of algorithm re-runs.
(a) Bellman-Ford
(b) Depth-first search
Figure 7: Pairwise mean distribution differences for DFS and BF for 20, 50, and 100 reruns for graphs of sizes 5 through 64.
Appendix C Validating the model output distributions
Going beyond our analysis of the validity of solutions obtained from the method presented in this paper, we investigate the coverage of the space of all solutions (i.e. the space of possible solutions we can obtain from executing an algorithm on a given graph $G$) of our methods. Using DFS as an example, we compare the average number of solutions when using DFS on graphs of a given size to the average number of unique and valid solutions when using our sampling method on the outputs of our model on the test data generated from those same graphs. This affords insight into the proportion of the space of possible solutions our method is capable of covering, which is particularly relevant to applications of Neural Algorithmic Reasoning in which using many different solutions is important, such as in cyberphysical systems.
C.1 Validating model output distributions for Bellman-Ford
(a) 5-vertex Graphs.
(b) 64-vertex Graphs.
Figure 8: The number of valid unique solutions found by our sampling methods on model output distributions and Bellman-Ford on the original graphs for 10 graphs and 25 samples.
Figure 8(a) shows that our sampling methods simply find the one valid solutions that exists for graphs of size 5 in all cases, which validates both our model and the sampling methods. In Figure 8(b), we can see that even for out-of-distribution graphs, the model output distributions are good enough for our sampling methods to return a high diversity of valid solutions, even within the margin error of straightforwardly applying BF to the input graph. This coverage of the solution space is very encouraging for further research into NAR with multiple solutions, as it shows the potential advantage of using NAR to find all possible solutions, which is relevant to many applications, as discussed in the body of the paper.
(a) 5-vertex Graphs.
(b) 64-vertex Graphs.
Figure 9: Mean edge reuse (the mean proportion of edges in common between resampled solutions for the same graph) for different Bellman-Ford sampling methods.
Figures 9 and 10 show that for both considered graph sizes, the proportions of edges in common among sampled solutions is similar across sampling methods (referred to further as mean edge reuse), which is unsurprising given the nature of single-source shortest paths, where only a certain subset of edges is relevant to finding the shortest paths.
(a) 5-vertex Graphs.
(b) 64-vertex Graphs.
Figure 10: Evolution of the mean edge reuse for different Bellman-Ford sampling methods.
C.2 Validating model output distributions for DFS
Figure 11(a) shows the marked advantage of AltUpwards over Upwards, as AltUpwards finds as many solutions as DFS on small graphs, where Upwards struggles to return any diversity of solutions. However, we can see the difficulty of emulating DFS using NAR on large graphs in Figure 11(b), where the diversity of valid solutions remains small over many samples from the model output distributions.
(a) 5-vertex Graphs.
(b) 64-vertex Graphs.
Figure 11: The number of valid unique solutions for different sampling methods for DFS for 10 graphs and 25 samples per graph.
(a) 5-vertex Graphs.
(b) 64-vertex Graphs.
Figure 12: Mean edge reuse for different sampling methods for DFS
For those solutions that are valid, however, Figures 12 and 13 show that our sampling methods yield solutions that are similarly diverse (as measured by the mean proportion of edges they share between them) as running DFS on the input graph. Thus, it seems that the quality of output distributions themselves ought to be improved, potentially with more targeted deep learning approaches, because the sampling methods seem to produce diverse solutions given good distributions.
(a) 5-vertex Graphs.
(b) 64-vertex Graphs.
Figure 13: Evolution of mean edge reuse for sampled DFS solutions for 10 graphs and 25 samples.
Appendix D Training Details
D.1 Model architecture and parameters
To give a good baseline, we use the default CLRS Benchmark NAR neural network [2]. Our processor network is an MPNN [10] with a two standard NAR adjustments: namely, gating and triplet reasoning [3]. Our encoder is standard: a linear layer initialised with xavier on scalars. Our decoder is 4 linear layers — three with hidden dimension (128) and a final layer with dimension one. The network is recurrent.
Parameter Value
Processor triplet_gmpnn
Encoder xavier_on_scalars
Decoder 4 linear layers
#Parameters 391225
#Hidden Units 128
Learning Rate 0.001
Batch Size 1
Training Graph Sizes 4, 7, 11, 13, 16
Training Set Size 1000
Testing Graph Sizes 5, 64
Testing Set Size 10
Table 3: Training parameters
D.2 Time characteristics
We run our experiments using a publicly available compute provider that uses NVIDIA P100 GPUs. Including evaluation of our model every 50 steps, training the NARs for 10,000 steps is very time-inexpensive. We summarize the training times in Table 4.
$n=5$ $n=64$
BF 240sec. 248sec.
DFS 531sec. 550sec.
Table 4: Training times for the conducted experiments in seconds